Dorie, 42 years old has a 12 year old nephew, in how many years will the aunt be twice as old as her nephew?

Answer by elima(1433)

Dorie, 42 years old has a 12 year old nephew, in how many years will the aunt be twice as old as her nephew?
You would use the same equation as #12057
42+x=2(12+x)
42+x=24+2x
Subtract x from both sides and 24 from both sides.
42-24=2x-x
18=x
check:
42+18=2(12+18)
60=2(30)
60=60
Hope this helps =)